Decide eventual extrema points to f(x) = (x – 3)ex in the interval [0,4]
Answer #1

Given function, f(x)=(x – 3)?e* Find the first derivative of this function. f(x) = (x – 3) e + 2(x – 3)e* =e* (x2 - 4x+3) F

f (1)=(1-3)? e = 4e f (3)=(3-3)? € = 0 f(0)=(0-3)?e° =9 f (4)=(4-3)?et = e* Hence, Local minima at x = 3 is 0. Local maxima